# STRAWBERRY: Finding haloes in the gravitational potential

### v1

@@
Tamara R. G. Richardson 11    Jens Stücker 22    Raul E. Angulo 1133
@@ 1 Introduction
Nonetheless, this halo definition has several drawbacks. Notably, spherical overdensities enforce spherical symmetry while haloes are typically ellipsoidal, and present many small-scale features small-scale such as subhaloes or caustics (e.g. Eisenstein & Loeb, 1995; Springel et al., 2004; Vera-Ciro et al., 2011; Despali et al., 2014; Bonnet et al., 2022). Moreover, spherical overdensities are not well suited to studying the evolution of haloes, due to the fact that the definition itself changes depending on the cosmological background, through an effect known as pseudo-evolution (Diemer et al., 2013). For example, this effect makes it more difficult to disentangle how much matter has been accreted over a given time period, as the spherical overdensity mass will increase over time even in the absence of accretion, simply because the background density is decreasing. This definition is also ill-suited to describe the abundance of haloes, as expressed through the halo mass function (HMF). Indeed, excursion-set theory predicts that the HMF should have no explicit cosmology dependence when expressed in the correct units (Press & Schechter, 1974; Bond et al., 1991), in other words, that the HMF should be universal. In practice, simulations have revealed this to not be the case (e.g. Tinker et al., 2008; Despali et al., 2016; Castro et al., 2021; Ondaro-Mallea et al., 2022). However, an open question remains, what is the origin of observed non universality? Is it generated through a physical process? or does it simply emerge from the mass definition?
@@
Figure 1: Slice of the boosted gravitational potential field centred on a halo, the dark green contours mark the saddle point energy, the bound population of particles is shown as magenta points.
@@
In $N$-body simulations, another popular choice of definition is the mass of Friends-of-Friends (FoF) groups, (e.g. Press & Davis, 1982; Springel et al., 2001; Roy et al., 2014), which depend on a linking length between individual particles that serves as a proxy for the local matter density. This has the advantage of removing the assumption of spherical symmetry, and being numerically efficient to implement. Moreover, this definition provides valuable insight as to the clumpy nature of haloes but comes at the expense of being more difficult to link to theory and not being directly applicable to observations. In addition, FoF has the tendency to detect chance alignment of particles, marking the presence of a halo where there is none, for example in Angulo et al. (2013) and Stücker et al. (2020, 2021) caustics and numerical fragments are detected by FoF as haloes in warm dark matter simulations. Furthermore, because FoF relies on the local density of particles, it may create unphysically large structures by bridging between haloes (see e.g. Knebe et al., 2011; Leroy et al., 2021). This particular issue stems from the fact that haloes lack a sharp boundary in terms of density.
@@
Nevertheless, in simulations the physical properties of dark matter, such as the density, velocity, and acceleration fields, are known. These allow to paint a more in depth picture of haloes by using more advanced definitions. While the edge of haloes is difficult to define from the density field, it is possible to do so in terms of dynamical properties. For instance, one can select particles based on their orbital status, considering only particles that have undergone their first apocentric passage (Adhikari et al., 2014; Diemer, 2017), or which have undergone shell crossing in three dimensions (Falck et al., 2012; Stücker et al., 2020) as part of the halo. Alternatively, one can define the boundary in energy space by introducing cuts in terms of kinetic energy (Garc´ıa et al., 2023; Salazar et al., 2024). These approaches are, however, much more computationally expensive, with the first methods requiring that one tracks the motion of every particle and the second requiring an internal fitting procedure to optimise the kinetic energy cut.
@@
that removes the locally-averaged large-scale gradient, $\left.\nabla_{\@vec{x}}\phi(\@vec{x})\right|_{\@vec{x}=\@vec{x}_{0}}$, an operation equivalent to studying a halo in a reference frame free-falling with an acceleration $\@vec{a}_{0}$. In this context, haloes can be defined as host potential wells, for which the boundary is set by a saddle point in the field which leads to a deeper potential well. Particles within the well that have energies below the saddle-point energy are bound to the structure, whereas particles with higher energies could escape. This definition of boundness can be seen as a generalization of commonly used self-binding checks (e.g. Springel et al., 2001; Behroozi et al., 2013) since it naturally includes the gravitational effect of surrounding structures. This is illustrated in Fig. 1, where we show a slice of the boosted gravitational potential field centred on a $10^{13}h^{-1}\mathrm{M}_{\odot}$ halo. Here we outline the saddle point energy, $\phi_{\rm sad}$, marking the dynamical edge of the halo, with a thick green line and show the particles which are bound to this potential well in magenta.
@@ 2.2 Potential Energy
$\nabla_{r}^{2}\phi_{r}$ $=4\pi G\rho_{r},$ (3)
where $\nabla_{r}$ denotes a derivative with respect to physical coordinates r\@vec{r} and $\rho_{r}$ is the physical density. Physical energy is given by
@@
where $\Lambda$ is the cosmological constant. The dark energy term is often omitted, as it is usually sufficiently small at the small distances that are of interest in physical space. However, we include it here for consistency with large scales.
In the comoving, frame we may write
@@
$\nabla_{x}^{2}\phi$ $=4\pi Ga^{2}(\rho_{r}-\rho_{\rm m})=\frac{3}{2}a^{-1}\Omega_{\rm m,0}H_{0}^{2}\delta,$ (8)
@@
where $x$ is the comoving coordinate, $a$ is the scale factor, v\@vec{v} is the canonical momentum which relates to peculiar velocities as $\@vec{v}_{\mathrm{pec}}=\@vec{v}/a$, $\rho_{\rm m}$ is the mean matter density of the universe, $H_{0}$ the Hubble parameter at $a=1$, $\delta$ the relative matter overdensity and $\phi$ is the peculiar potential. Note that conventions on the definition of the peculiar potential may differ by a factor $a$, for example the cosmological simulation code Gadget (Springel et al., 2021a) outputs $\phi_{\rm gadget}=a\cdot\phi$ in snapshots. Importantly the physical potential and the peculiar potential differ by a quadratic term
@@
where the third line is assuming a flat universe without radiation $\Omega_{\Lambda}=1-\Omega_{\rm m}$. We can calculate $E$ from the two different frames and get consistent results for the same particle. However, the split between which part of the energy is considered as potential energy and kinetic energy is frame-dependent. For a binding check, this implies that we may define a procedure which is consistent between resting and expanding reference frames by requiring $E<E_{\infty}$ where $E_{\infty}$ is some reference energy above which a particle may escape the potential well. The motivation of using the physical energy $E$ is that it is closest to a conserved quantity for static physical systems that have decoupled from the expansion of the Universe.
@@
$\partial_{r}\phi_{r}$ $=\frac{4\pi G}{3}r\rho_{\rm m}[1+\overline{\delta}(r)]>0,$ (12)
we find that it is monotonically increasing, and, as such, has no saddle point. In opposition to the peculiar potential,
@@
which reaches $\partial_{x}\phi\rightarrow 0$ only at infinity. In practice, neither case is convenient.
@@
We note that the choice of the potential that is used to define the saddle-point energy level leaves a degree of freedom that may seem somewhat arbitrary. However, in practice the difference between the different potential notions is almost always negligible, since they only tend to differ at much larger radii than the location of the saddle-point. We will show this in a bit more detail in App. A.3.
@@ 2.3 The boosted potential
Definitions of energy depend on the adopted reference frame. For example, a Galilei transformation changes the kinetic energy term and it modifies the time-dependence of the potential. Furthermore, a transformation into an accelerated reference system with an arbitrary time-dependent offset $\@vec{x}_{0}(t)$,
@@
with $\@vec{v}_{0}=a^{-2}\dot{\@vec{x}}_{0}$, introduces an additional fictitious force in the transformed coordinates:
@@
with $a_{0}(t)=\dot{\@vec{v}}_{0}(t)$. With the equation of motion $\dot{v}=-\nabla_{x}\phi$, we may interpret this as a modification of the potential
@@
Since the potential arises from the mass distribution, the explicit time-dependence of the mass distribution, $\rho(\@vec{x}^{\prime},t)$, should be small in an optimal reference frame. As such, given a set of particles $\{\@vec{x}_{i},\@vec{v}_{i}\}$ – e.g. all particles in a region of interest – we choose our reference frame so that the centre of mass, the centre of mass velocity, and the centre of mass acceleration are fixed at 0.
$\langle\@vec{x}^{\prime}\rangle(t)$ $=0,$ (23)
$\langle\@vec{v}^{\prime}\rangle(t)$ $=0,$ (24)
@@
in the accelerated reference frame. This potential is the field that will be used in the following section to define the dynamical boundary of haloes, and energies of particles by replacing $\phi_{*}$ with $\phi_{*,b}$ in Eqs. (17) and (18). Indeed, this definition grants us all of the benefits discussed above by both minimising the explicit time dependence of the potential while also guaranteeing that there will always exist a saddle point to which we can define the boundary of the potential well.
@@ 3.1 Moving about the simulation
3.2 Initialization and reference frame
@@ 4.1 When and where do particles become bound to a halo?
Asking ourselves the question, ‘when and where do particles become bound to haloes?’ echoes dynamical halo boundary definitions, where the boundary of haloes is set not in terms of density but the dynamics of particles such as particle orbits or energies (Diemer, 2017; Garc´ıa et al., 2023; Salazar et al., 2024). To answer this question, we select a sample of 100 haloes with bound masses between $5\cdot 10^{13}\ h^{-1}\mathrm{M}_{\odot}$ and $1.5\cdot 10^{14}\ h^{-1}\mathrm{M}_{\odot}$ for which we track the assembly history from $a=0.5$ to $a=1$, and simultaneously track all particles which enter within a spherical region of radius $R<1.5\sqrt{3}\simeq 2.6\ h^{-1}{\rm Mpc}$ centred on the position of the main progenitors. During the tracking procedure we travel backwards in time defining the main progenitor group as the FoF halo within the tracked region which contains the most particles which will be bound to the main descendant halo in the following snapshot. Once the main progenitor is detected, we recentre the region on its location before proceeding with the particle assignment procedure described in Sect. 3. Within each region we track the individual positions, velocities, boosted potential values, and energies of all particles.
@@
From the resulting dataset, we estimate the energy histories of all particles that enter theses 100 regions. Among these energy tracks, we select three subsamples of particles which become bound to their host haloes at scale factors $a\in[0.6;0.7]$, $a\in[0.7,0.8]$, and $a\in[0.8,0.9]$. In order to stack energy histories originating from different haloes, we normalise the individual histories by the persistence of their host halo at $a=1$. In Fig. 4 we show the evolution of the median energies of all three subsamples along with the evolution of the median persistence, normalised at $a=1$. Over time, the energy of particles continuously grows slowly by about $20-30\%$ between $a=0.5$ and $a=1.0$. On the other hand, the persistence grows at a significantly faster rate, almost doubling since $a=0.5$. Since particles become bound when they cross the persistence threshold, the core reason that particles become bound here is that the persistence grows – primarily by additional mass being accreted outside of the orbital radius of the particles of interest.
Note that this is quite different than what would be observed if the self-potential was used to define energies. A notable difference is that the self-potential is normalized to zero at infinity, whereas our potential notion is normalized to be zero at the minimum. In the self-potential, particles become bound because they decrease their self-energy and cross the zero-energy level. E.g. if material is accreted at far larger radii than the orbital radius of a particle of interest, then the self-energy decreases, whereas the energy defined here would be unaffected (while the persistence of the potential increases). We argue that this is a desirable property and suggest that normalizing energies to zero at the minimum may make comparisons of energy levels across time more meaningful. Note that such a perspective also simplifies understanding other processes, like tidal stripping (e.g. Drakos et al., 2020; Errani et al., 2022; Stücker et al., 2023).
To gauge at what point of their orbit particles become bound, in Fig. 5 we compute the fraction of bound particles after a given number of dynamical times (Jiang & van den Bosch, 2016),
@@
We find that the fraction of bound particles begins to increase as soon as particles enter the the saddle point contour of the halo, we mark this moment with a green circle, with the error bars around this point marking the standard deviation of $N_{\rm dyn}$ between the particle entering the halo and its first pericentric passage. Most particles that become bound at this stage are particles that are accreted with lower-than-average energies. As we approach the pericentre, the fraction of bound particles increases rapidly, with 40% of particles being bound to the structure at the first pericentric passage, marked by an orange star [^3] This point is by definition at $N_{\rm dyn}=0$ and hence does not have any dispersion.. At the first apocentre, marked by a red square, we find that 75% of particles are considered bound to the structure. Finally, we observe that after two dynamical times, 90% of the particles are considered bound. From these considerations, we confirm that gravitational binding occurs on a relatively short timescale, typically only a few dynamical times. These observations support the motivations behind particle selection criteria based on orbital status. For instance, one may select particles based on whether it has passed its first pericentre or apocentre (e.g Adhikari et al., 2014; Diemer, 2017), meaning parallels can be drawn between these types of mass definition and the one presented in this work.
@@
The binding of particles can further be explored in phase space. Indeed, it is generally accepted that in phase space, CDM haloes can be represented as ravelled three-dimensional manifolds (Shandarin & Zeldovich, 1989). When projected in terms of radius, $r$, and radial velocity, $v_{\rm r}$, haloes appear as a spiral pattern, with the most recently accreted material being deposited on the outskirts of the halo and older material being concentrated in the centre. The halo is then connected to the background medium through a stream of infalling material (see e.g. Wang et al., 2011; Adhikari et al., 2014; Zavala & Frenk, 2019).
In Fig. 6, we show the stacked radial phase space density of the same 100 haloes, The phase space density is split into bound (magenta), unbound and infalling (green), and unbound and splashback (orange) populations. We select bound particles according to the boosted potential criterion presented above and separate the unbound sets using the orbits of particles. The brightness of each contour is set to enclose 50%, 84%, and 95%, respectively from darkest to lightest, of the total population, including both bound and unbound population. Here, the bound population occupies a triangular region around the centre of the halo. This population is relatively symmetric in terms of radial velocity. As such it is reasonable to think that this population is virialised, we discuss this quantitatively in Sect. 4.3. This region of phase space, is flanked at larger radii and radial velocities by material classified as unbound. We observe that the second component can be split into two parts, One with negative radial velocities, which we refer to as the first infall stream and a second with positive radial velocities, which we refer to as the first splashback stream. We note that the first splashback stream is less dens than the first infall stream. This is in line with our previous observation, where we find that by the first pericentric passage approximately half of particles are already bound to the halo, as such they will not be included here. This configuration of bound and unbound particles is consistent with the kinetic energy - radius cuts proposed by Salazar et al. (2024) to filter out recently accreted material. Indeed, when plotted in terms of kinetic energy and radius, the distinction between bound and unbound particles in Fig. 6, follow the valley features used in Salazar et al. (2024) to define the orbiting and infalling populations.

### v2

@@
Tamara R. G. Richardson    Jens Stücker    Raul E. Angulo
@@ 1 Introduction
Nonetheless, this halo definition has several drawbacks. Notably, it enforces spherical symmetry while haloes are typically ellipsoidal, and present many small-scale features such as subhaloes or caustics (e.g. Eisenstein and Loeb, 1995; Springel et al., 2004; Vera-Ciro et al., 2011; Despali et al., 2014; Bonnet et al., 2022). Moreover, spherical overdensities are not well suited to studying the evolution of haloes, due to the fact that the definition itself changes depending on the cosmological background, through an effect known as pseudo-evolution (Diemer et al., 2013). For example, this effect makes it more difficult to disentangle how much matter has been accreted over a given time period, as the spherical overdensity mass will increase over time even in the absence of accretion, simply because the background density is decreasing. This definition is also ill-suited to describe the abundance of haloes, as expressed through the halo mass function (HMF). Indeed, excursion-set theory predicts that the HMF should have no explicit cosmology dependence when expressed in the correct units (Press and Schechter, 1974; Bond et al., 1991), in other words, that the HMF should be universal. In practice, simulations have revealed this to not be the case (e.g. Tinker et al., 2008; Despali et al., 2016; Castro et al., 2021; Ondaro-Mallea et al., 2022). However, an open question remains, what is the origin of observed non universality? Is it generated through a physical process? or does it simply emerge from the mass definition?
@@
Figure 1: Slice of the boosted gravitational potential field centred on a halo, the dark green contours mark the saddle point energy, $\phi_{\rm sad}$, the bound population of particles is shown as magenta points.
@@
In $N$-body simulations, another popular choice of definition is the mass of Friends-of-Friends (FoF) groups, (e.g. Press and Davis, 1982; Springel et al., 2001; Roy et al., 2014), which depend on a linking length between individual particles that serves as a proxy for the local matter density. This has the advantage of removing the assumption of spherical symmetry, and being numerically efficient to implement. Moreover, this definition provides valuable insight as to the clumpy nature of haloes but comes at the expense of being more difficult to link to theory and not being directly applicable to observations. In addition, FoF has the tendency to detect chance alignment of particles, marking the presence of a halo where there is none, for example in Angulo et al. (2013) and Stücker et al. (2020, 2021) caustics and numerical fragments are detected by FoF as haloes in warm dark matter simulations. Furthermore, because FoF relies on the local density of particles, it may create unphysically large structures by bridging between haloes (see e.g. Knebe et al., 2011; Leroy et al., 2021). This particular issue stems from the fact that haloes lack a sharp boundary in terms of density.
@@
Nevertheless, in simulations the physical properties of dark matter, such as the density, velocity, and acceleration fields, are known. These allow to paint a more in depth picture of haloes by using more advanced definitions. While the edge of haloes is difficult to define from the density field alone, (see e.g. Eisenstein and Hut, 1998; Springel et al., 2001; Aubert et al., 2004; Codis et al., 2015, 2018), it is possible to do so in terms of dynamical properties. For instance, one can select particles based on their orbital status, considering only particles that have undergone their first apocentric passage (Adhikari et al., 2014; Diemer, 2017), or which have undergone shell crossing in three dimensions (Falck et al., 2012; Stücker et al., 2020) as part of the halo. Alternatively, one can define the boundary in energy space by introducing cuts in terms of kinetic energy (García et al., 2023; Salazar et al., 2025). These approaches are, however, much more computationally expensive, with the first methods requiring that one tracks the motion of every particle and the second requiring an internal fitting procedure to optimise the kinetic energy cut.
@@
that removes the locally-averaged large-scale gradient, $\left.\nabla_{\@vec{x}}\phi(\@vec{x})\right|_{\@vec{x}=\@vec{x}_{0}}$, an operation equivalent to studying a halo in a reference frame free-falling with an acceleration $\@vec{a}_{0}$. With the equivalence principle implying that this type of transformation does not impact the internal dynamics of the system. In this context, haloes can be defined as host potential wells, for which the boundary is set by a saddle point in the field which leads to a deeper potential well. Particles within the well that have energies below the saddle-point energy are bound to the structure, whereas particles with higher energies could escape. This definition of boundness can be seen as a generalization of commonly used self-binding checks, (e.g. Springel et al., 2001; Behroozi et al., 2013) since it naturally includes the gravitational effect of surrounding structures. This is illustrated in Fig. 1, where we show a slice of the boosted gravitational potential field centred on a $10^{13}h^{-1}\mathrm{M}_{\odot}$ halo. Here we outline the saddle point energy, $\phi_{\rm sad}$, marking the dynamical edge of the halo, with a thick green line and show the particles which are bound to this potential well in magenta.
@@ 2.2 Potential Energy
$\nabla_{r}^{2}\phi_{r}$ $=4\pi G\rho(\@vec{r}),$ (3)
where $\nabla_{r}$ denotes a derivative with respect to physical coordinates r\@vec{r} and $\rho(\@vec{r})$ is the physical density. Physical energy is given by
@@
where $\Lambda$ is the cosmological constant. The dark energy term is often omitted, as it is usually sufficiently small at the small distances that are of interest in physical space. However, since it has an effect on larger scales, it is included for consistency.
In the comoving frame, we may write
@@
$\nabla_{x}^{2}\phi$ $=4\pi Ga^{2}[\rho(\@vec{r})-\rho_{\rm m}]=\frac{3}{2}a^{-1}\Omega_{\rm m,0}H_{0}^{2}\delta,$ (8)
@@
where $x$ is the comoving coordinate, $a$ is the scale factor, v\@vec{v} is the canonical momentum which relates to peculiar velocities as $\@vec{v}_{\mathrm{pec}}=\@vec{v}/a$, $\rho_{\rm m}$ is the mean matter density of the universe, $H_{0}$ the Hubble parameter at $a=1$, $\delta$ the relative matter overdensity, $\phi$ is the peculiar potential, and the dot marks derivatives with respect to the proper time for a comoving observer. Note that conventions on the definition of the peculiar potential may differ by a factor $a$, for example the cosmological simulation code Gadget (Springel et al., 2021b) outputs $\phi_{\rm gadget}=a\cdot\phi$ in snapshots. Importantly, the physical potential and the peculiar potential differ by a quadratic term
@@
where the third line is assuming a flat universe without radiation, so $\Omega_{\Lambda}=1-\Omega_{\rm m}$. We can calculate $E$ from the two different frames and get consistent results for the same particle. However, the split between which part of the energy is considered as potential energy and kinetic energy is frame-dependent. For a binding check, this implies that we may define a procedure which is consistent between resting and expanding reference frames by requiring $E<E_{\infty}$ where $E_{\infty}$ is some reference energy above which a particle may escape the potential well. The motivation of using the physical energy $E$ is that it is closest to a conserved quantity for static physical systems that have decoupled from the expansion of the Universe.
@@
$\partial_{r}\phi_{r}$ $=\frac{4\pi G}{3}r\rho_{\rm m}[1+\overline{\delta}(r)],$ (12)
we find that it is monotonically increasing ($\partial_{r}\phi_{r}$ ¿ 0), and, as such, has no saddle point. In contrast, for the peculiar potential, it is
@@
which reaches $\partial_{r}\phi\rightarrow 0$ only at infinity. In practice, neither case is convenient.
@@
We note that the choice of the potential that is used to define the saddle-point energy level leaves a degree of freedom that may seem somewhat arbitrary. However, in practice the difference between the different potential notions is almost always negligible, since they only tend to differ at much larger radii than the location of the saddle-point. We will show this in more detail in App. A.3.
@@ 2.3 The boosted potential
Definitions of energy depend on the adopted reference frame. For example, a Galilean transformation changes the kinetic energy term and it modifies the time-dependence of the potential. Furthermore, a transformation into an accelerated reference system with an arbitrary time-dependent offset $\@vec{x}_{0}(t)$,
@@
with $\@vec{v}_{0}=a^{2}\dot{\@vec{x}}_{0}$, introduces an additional fictitious force in the transformed coordinates:
@@
with $\@vec{a}_{0}(t)=\dot{\@vec{v}}_{0}(t)$. With the equation of motion $\dot{\@vec{v}}=-\nabla_{x}\phi$, we may interpret this as a modification of the potential,
@@
Since the potential arises from the mass distribution, the explicit time-dependence of the mass distribution, $\rho(\@vec{x}^{\prime},t)$, should be small in an optimal reference frame. As such, given a set of particles $\{\@vec{x}_{i},\@vec{v}_{i}\}$ – e.g. all particles in a region of interest – we choose our reference frame so that the centre of mass, the centre of mass velocity, and the centre of mass acceleration are fixed at 0:
$\langle\@vec{x}^{\prime}\rangle(t)$ $=0;$ (23)
$\langle\@vec{v}^{\prime}\rangle(t)$ $=0;$ (24)
@@
in the accelerated reference frame. This potential is the field that will be used in the following section to define the dynamical boundary of haloes, and energies of particles by replacing $\phi_{*}$ with $\phi_{\rm b,*}$ in Eqs. (17) and (18). Indeed, this definition grants us all of the benefits discussed above by both minimising the explicit time dependence of the potential while also guaranteeing that there will always exist a saddle point to which we can define the boundary of the potential well.
@@ 3.1 Moving about the simulation
3.2 Initialisation and reference frame
@@ 4.1 When and where do particles become bound to a halo?
Asking ourselves the question, ‘when and where do particles become bound to haloes?’ echoes dynamical halo boundary definitions, where the boundary of haloes is set not in terms of density but the dynamics of particles such as particle orbits or energies (Diemer, 2017; García et al., 2023; Salazar et al., 2025). To answer this question, we select a sample of 100 haloes with bound masses between $5\cdot 10^{13}\ h^{-1}\mathrm{M}_{\odot}$ and $1.5\cdot 10^{14}\ h^{-1}\mathrm{M}_{\odot}$ for which we track the assembly history from $a=0.5$ to $a=1$, and simultaneously track all particles which enter within a spherical region of radius $R<1.5\sqrt{3}\simeq 2.6\ h^{-1}{\rm Mpc}$ centred on the position of the main progenitors. During the tracking procedure we travel backwards in time defining the main progenitor group as the FoF halo within the tracked region which contains the most particles which will be bound to the main descendant halo in the following snapshot. Once the main progenitor is detected, we recentre the region on its location before proceeding with the particle assignment procedure described in Sect. 3. Within each region we track the individual positions, velocities, boosted potential values, and energies of all particles.
@@
From the resulting dataset, we estimate the energy histories of all particles that enter theses 100 regions. Among these energy tracks, we select three subsamples of particles which become bound to their host haloes at scale factors $a\in[0.6;0.7]$, $a\in[0.7,0.8]$, and $a\in[0.8,0.9]$. In order to stack energy histories originating from different haloes, we normalise the individual histories by the persistence of their host halo at $a=1$. In Fig. 4 we show the evolution of the median energies of all three subsamples along with the evolution of the median persistence, normalised at $a=1$. Over time, the energy of particles continuously grows slowly by about $20\%$ to $30\%$ between $a=0.5$ and $a=1.0$ up to the time when they are accreted, after which their energy stays constant. On the other hand, the persistence grows at a significantly faster rate, almost doubling since $a=0.5$. Since particles become bound when they cross the persistence threshold, the core reason that particles become bound here is that the persistence grows – primarily by additional mass being accreted outside of the orbital radius of the particles of interest.
Note that this is quite different than what would be observed if the self-potential was used to define energies. A notable difference is that the self-potential is normalized to zero at infinity, whereas our potential notion is normalized to be zero at the minimum. In the self-potential, particles become bound because they decrease their self-energy and cross the zero-energy level. E.g. if material is accreted at far larger radii than the orbital radius of a particle of interest, then the self-energy decreases, whereas the energy defined here would be unaffected (while the persistence of the potential increases). This can be seen in Fig. 4 where the energy history of particles flattens out after they become bound. We argue that this is a desirable property, and suggest that normalizing energies to zero at the minimum may make comparisons of energy levels across time more meaningful. Indeed, we can see that the energies remain clearly separated and maintain their ordering after accretion, indicating that once the halo is virialised it maintains a memory of the accretion process. Note that such a perspective also simplifies understanding other processes, like tidal stripping (e.g. Drakos et al., 2020; Errani et al., 2022; Stücker et al., 2023).
To gauge at what point of their orbit particles become bound, in Fig. 5 we compute the fraction of bound particles after a given number of dynamical times (Jiang and van den Bosch, 2016),
@@
We find that the fraction of bound particles begins to increase as soon as particles enter the saddle point contour of the halo, we mark this moment with a green circle, with the error bars around this point marking the standard deviation of $N_{\rm dyn}$ between the particle entering the halo and its first pericentric passage. Most particles that become bound at this stage are particles that are accreted with lower-than-average energies. As we approach the pericentre, the fraction of bound particles increases rapidly, with 50% of particles being bound to the structure at the first pericentric passage, marked by an orange star [^3] This point is by definition at $N_{\rm dyn}=0$ and hence does not have any dispersion.. At the first apocentre, marked by a red square, we find that 75% of particles are considered bound to the structure. Finally, we observe that after two dynamical times, 90% of the particles are considered bound. From these considerations, we confirm that gravitational binding occurs on a relatively short timescale, typically only a few dynamical times. These observations support the motivations behind particle selection criteria based on orbital status. For instance, one may select particles based on whether it has passed its first pericentre or apocentre (e.g Adhikari et al., 2014; Diemer, 2017), meaning parallels can be drawn between these types of mass definition and the one presented in this work.
@@
The binding of particles can further be explored in phase space. Indeed, it is generally accepted that in phase space, CDM haloes can be represented as ravelled three-dimensional manifolds (Shandarin and Zeldovich, 1989). When projected in terms of radius, $r$, and radial velocity, $v_{\rm r}$, haloes appear as a spiral pattern, with the most recently accreted material being deposited on the outskirts of the halo and older material being concentrated in the centre. The halo is then connected to the background medium through a stream of infalling material (see e.g. Wang et al., 2011; Adhikari et al., 2014; Zavala and Frenk, 2019).
In Fig. 6, we show the stacked radial phase space density of the same 100 haloes, The phase space density is split into bound (magenta), unbound and infalling (green), and unbound and splashback (orange) populations. We select bound particles according to the boosted potential criterion presented above and separate the unbound sets using the orbits of particles. The brightness of each contour is set to enclose 50%, 84%, and 95%, respectively from darkest to lightest, of the total population, including both bound and unbound population. Here, the bound population occupies a triangular region around the centre of the halo. This population is relatively symmetric in terms of radial velocity. As such it is reasonable to think that this population is virialised, we discuss this quantitatively in Sect. 4.3. This region of phase space, is flanked at larger radii and radial velocities by material classified as unbound. We observe that the second component can be split into two parts, one with negative radial velocities, which we refer to as the first infall stream and a second with positive radial velocities, which we refer to as the first splashback stream. We note that the first splashback stream is less dense than the first infall stream. This is in line with our previous observation, where we find that by the first pericentric passage approximately half of particles are already bound to the halo, as such they will not be included here. This configuration of bound and unbound particles is consistent with the kinetic energy - radius cuts proposed by Salazar et al. (2025) to filter out recently accreted material. Indeed, when plotted in terms of kinetic energy and radius, the distinction between bound and unbound particles in Fig. 6, follow the valley features used in Salazar et al. (2025) to define the orbiting and infalling populations.